OpenStudy (anonymous):

Find four consecutive odd integers if the product of the two smaller integers is 112 less than the product of the two larger integers. @alli14344 @mathstudent55 @matty546 @jiraya

OpenStudy (anonymous):

We began by assuming that the 4 odd integers are consecutive, and everything worked just fine. So, although there may well be more than one correct answer to the question as it's stated, we have a hunch they're supposed to be consecutive, and that detail was accidentally left out of the question. Call the 4 consecutive odd integers (2x - 3), (2x - 1), (2x + 1), and (2x + 3). Product of the two larger ones = (4x2 + 8x + 3) Product of the two smaller ones = (4x2 - 8x + 3) Smaller product + 112 = larger product . . . . . (4x2 - 8x + 3) + 112 = (4x2 + 8x + 3) Eliminate parentheses . . . . . 4x2 - 8x + 115 = 4x2 + 8x + 3 Subtract (4x2 + 3) from each side . . . . . -8x + 112 = 8x Add 8x to each side . . . . . 112 = 16x Divide each side by 16 . . . . . x = 7 The consecutive odd integers are: 11, 13, 15, and 17. Check: 11 x 13 = 143 15 x 17 = 255 255 - 143 = 112 Yippee!

OpenStudy (anonymous):

ok thanks